OpenStudy (anonymous):

What is 3/4 plus 5/6?

OpenStudy (anonymous):

convert them both to a common denominator and add

OpenStudy (anonymous):

9/12 + 10/12 = 19/12

OpenStudy (anonymous):

It is 1unit and 7/12

OpenStudy (anonymous):

yes

OpenStudy (lilg132):

wolframalpha.com

OpenStudy (anonymous):

are you computing these or just guessing from your multiple choice?

OpenStudy (anonymous):

I deserve a medel just kidding no i really mean it or do I no its your decision to give me a medel or not

OpenStudy (anonymous):

Do you think I need a medel

OpenStudy (lilg132):

no

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

38/24=19/12